OpenStudy (anonymous):

you deposit 500 in an ccount that pays 4% annual interest.how long does it take the balance to reach the following amounts 1000 when interest is compouded compounded quaterly. A)4.4 b)17.4 c)17.7 D)69.7

OpenStudy (radar):

1000=500(1+.04/4)^4t 1000=500(1.01)^4t 2=1.01^4t 4tlog 1.01=log2

OpenStudy (radar):

4t(.0043213)=.30103 .0172855t=.30103 t=17.415 years

OpenStudy (radar):

The formula for compound interest: \[A=P(1 + r/n)^{nt}\]Where A is the Amount, P is the Principal, r is the rate, n is the compounding number per year. and t is the number of years.

OpenStudy (radar):

This problem the t was the unknown.